a stone is thrown straight upward with a speed of 20m/s. it is caught on its way down at a point 5.0m above where it was thrown


1
Expert's answer
2021-03-18T14:04:18-0400
"v^2=u^2+2gh\\v^2=20^2+2(9.8)(5)\\v=22.3\frac{m}{s}"